OpenStudy (chelseatheruler):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3)

OpenStudy (jdoe0001):

what would you say is the slope of -> y = 2x + 2 ?

OpenStudy (chelseatheruler):

2

OpenStudy (jdoe0001):

ok, perpendicular lines have a NEGATIVE RECIPROCAL slopes what does that mean? well, if this has a slope of 2 any perpendicular one to it will have reciprocal of 2 1/2 negativize that -1/2

OpenStudy (jdoe0001):

so, now you have the slope of the other line, and a point it passes through so use both in the point-slope form $$ \large { y-y_1=m(x-x_1)}$$

OpenStudy (chelseatheruler):

okay. I dont understand how to do that.

OpenStudy (jdoe0001):

alright, so, we know the slope is -1/2, we know that a point it passes through is (6, 3) so, using \(y-y_1=m(x-x_1)\) we'll have $$ (\color{red}{6}, \color{red}{3}), m = \color{blue}{-\frac{1}{2}}\\ y-(\color{red}{3})=\color{blue}{-\frac{1}{2}}(x-(\color{red}{6})) $$

OpenStudy (jdoe0001):

now just solve for "y", and you'd get the line equation